OpenStudy (anonymous):

What is 4x * 7y / 24^2 when X = 24 and Y = 2? If you help I will fan you and write a testimonial for you and I will give a medal to the guy or girl that has the best answer!

OpenStudy (anonymous):

You just put a 24 wherever you see an x, hence x = 24. and a 2 wherever you see a y, hence y = 2 =4(x) * 7(y) /(24^2) =4(24) * 7(2) / (24^2) =(96 * 14) / 576 =1344/576 2 1/3